### v1

@@ 3 Vocabulary Parallelism in Pipeline Parallelism
Figure 4: Computation graph of the output layer after partitioning across the vocabulary dimension. There are three all-reduce communications across all devices.
@@ 4.1 A Naïve Approach
The partitioned output layer can be grouped into three computation passes $\textit{F}_{1}$, $\textit{F}_{2}$ and B, separated by three all-reduce communications involving the maximum of logits, the sum of logits and the input gradients respectively. We can overlap these all-reduce communications with transformer layer computation by placing them in a separate stream, as shown in Figure 5.
@@ 4.4 Backward Phase Optimization
We also observe that all three all-reduce communications can be done after computing the matrix multiplications for the input gradients. Note that
@@ 6.1 Implementation
In practice, scheduling under the assumption that backward takes twice time as forward might introduce unnecessary bubbles, especially when these values differ significantly. As a result, we profile the run time of the passes and schedule the $S$ and $T$ passes accordingly.
@@
In all our experiments, we untied the input and output embedding weights. However, note that our method eases embedding weight tying since the input and output layers have identical device placement.
@@ Appendix D V-Half Pipeline Scheduling
Appendix E Detailed Experiment Data

### v2

@@ 3 Vocabulary Parallelism in Pipeline Parallelism
Figure 4: Computation graph of the output layer after partitioning across the vocabulary dimension. There are three all-reduce / reduce communications across all devices.
@@ 4.1 A Naïve Approach
The partitioned output layer can be grouped into three computation passes $\textit{F}_{1}$, $\textit{F}_{2}$ and B, separated by three all-reduce / reduce communications involving the maximum of logits, the sum of logits and the input gradients respectively. We can overlap these all-reduce communications with transformer layer computation by placing them in a separate stream, as shown in Figure 5.
@@ 4.4 Backward Phase Optimization
We also observe that all three all-reduce / reduce communications can be done after computing the matrix multiplications for the input gradients. Note that
@@ 6.1 Implementation
Scheduling under the assumption that backward takes twice time as forward might introduce unnecessary bubbles, especially when these values differ significantly. As a result, we profile the run time of the passes and schedule the $S$ and $T$ passes accordingly [^1] The profiling verifies whether the backward pass runs approximately twice the time as the forward pass (unrelated to the vocabulary layer), in order to insert the vocabulary passes appropriately. The pipeline schedule generated will remain unchanged unless this ratio deviates by a certain threshold. In practice, we find that the difference is negligible in most transformer networks, and these differences would not change the pipeline schedule. Hence, this profiling could be viewed as optional..
@@
Note that our method makes tying input and output embedding weights easier as the input and output embedding weights now have the same device placement and can use the shared weight tensor. This saves GPU memory and avoids the additional all-reduce to synchronize gradients. However, in all our experiments, we adapted the more difficult setting, untying the input and output embedding weights, since it is adapted by some open source models like Llama 3 (Dubey et al., 2024).
@@ 6.2 Setup
Our experiments use pure pipeline parallelism to verify that our method improves pipeline parallelism as expected. Given that our method is orthogonal to tensor and data parallelism, the conclusions can be generalized to the production environment where pipeline parallelism is used together with tensor and data parallelism.
@@ Appendix D V-Half Pipeline Scheduling
Appendix E Correctness Evaluation
Our implementation is based on the open-source Megatron-LM project Narayanan et al. (2021). We compare the convergence curves of our implementation to that of the original Megatron-LM codebase to verify our implementation’s correctness. The configurations follow the 4B model in section 6.2 with a vocabulary size of 256K, trained with 8 GPUs. Additionally, we verify that our implementation also works correctly with tensor parallelism, by using a pipeline parallel size of 4 and a tensor parallel size of 2.
The convergence curves are shown in Figure 17. It is shown that our implementation maintains correctness, albeit with some small numerical differences.
Figure 17: Convergence curves of our implementation against the original Megatron-LM codebase
Appendix F Detailed Experiment Data
@@
Appendix G Artifact Appendix
G.1 Abstract
This section will outline the setup and experimental workflow of Balancing Pipeline Parallelism with Vocabulary Parallelism conducted on a single server equipped with 8 A100 GPUs.
G.2 Artifact check-list (meta-information)
  • Algorithm: Vocabulary Parallelism, Online Softmax
  • Program: Not used
  • Compilation: Nvidia nvcc version 12.4, already in the container
  • Transformations: Not used
  • Binary: will be compiled on a target platform
  • Data set: Customized C4 hosted in Huggingface
  • Binary: will be compiled on a target platform.
  • Run-time environment: Ubuntu 20.04.6. Needs docker. Requires root.
  • Hardware: Server with 8 Nvidia A100 GPUs 80GB HBM.
  • Run-time state: Server is idle.
  • Execution: Takes at least 4 hours to complete.
  • Metrics: Peak Memory, MFU
  • Output: The data printed on console. The output of Quick Experiment is the key result.
  • Experiments: Elaborated in the Installation and Experiment workflow sections.
  • How much disk space required (approximately)?: 30 GB
  • How much time is needed to prepare workflow (approximately)?: 30 minutes
  • How much time is needed to complete experiments (approximately)?: 4 hours
  • Publicly available?: Yes
  • Code licenses (if publicly available)?: Apache License
  • Data licenses (if publicly available)?: odc-by
  • Workflow framework used?: No
  • Archived (provide DOI)?:
G.3 Description
This experiment consists of 2 parts:
  • Quick Experiment to quickly verify the result on a specific case.
  • Full Experiment to run all cases on an 8-GPU server.
The Full Experiment employs the settings in table 7 same as the paper. The Quick Experiment focuses on a specific case where the sequence length is 4096 and the vocabulary size is 256k. We use Megatron-LM on which all methods are implemented to run training benchmarks.
Table 7: Artifact Settings used in experiments on 1F1B schedule.
Pipelines (GPUs) 8
Model Size $\approx$ 4B
Layers 32
Attention Heads 24
Hidden Size 3072
Sequence Length 2048 / 4096
Microbatch Size 1
Number of Microbatches 128
Vocabulary Size 32k / 64k / 128k / 256k
G.3.1 How delivered
The code locates on https://github.com/sail-sg/VocabularyParallelism. The dataset will be automatically downloaded in experiment scripts.
G.3.2 Hardware dependencies
The tests should be conducted in a server with 8 Nvidia A100 GPUs, 80GB HBM.
G.3.3 Software dependencies
  • CUDA Driver Version: 535.54.03
  • CUDA Version 12.2
  • Docker
  • NVIDIA Container Toolkit
G.3.4 Data sets
The dataset is customized from C4 hosted in Huggingface
https://huggingface.co/datasets/mtyeung/vocab_parallel_sample_dataset supporting various sequence length. It will be automatically downloaded in experiment scripts.
G.4 Installation
Run a container:
docker run --name vocab_torch24 \
--network=host -d \
--runtime=nvidia --gpus all \
--ipc=host --ulimit memlock=-1 --ulimit stack=67108864 \
--privileged=true \
nvcr.io/nvidia/pytorch:24.03-py3 sleep infinity
Get inside the container, and clone the codes:
# Enter the container
docker exec -it vocab_torch24 bash
# Clone the codes
git clone https://github.com/sail-sg/VocabularyParallelism.git
cd VocabularyParallelism
Note that all the following commands should be run inside the VocabularyParallelism directory.
G.5 Experiment workflow
G.5.1 Quick Experiment
The quick experiment runs all the methods (baseline, redis, interlaced, vocab-1, vocab-2) on a specific setting in the paper:
  • Sequence Length: 4096
  • Vocabulary Size: 256k
The experiment will show 2 key results:
  • Peak Memory
  • MFU
Run all the methods one by one:
bash artifact/quick_exp.sh run baseline
bash artifact/quick_exp.sh run redis
bash artifact/quick_exp.sh run interlaced
bash artifact/quick_exp.sh run vocab-1
bash artifact/quick_exp.sh run vocab-2
This will automatically download the dataset from huggingface and run the training experiments. The log containing the result will locate in quick-logs/<method>/stdout.log. Each method should take around 6 minutes to complete.
Then run this to collect the results:
bash artifact/quick_exp.sh show-result
G.5.2 Full Experiment
This will run all experiments on a single server with 8 A100 GPUs.
The whole experiment will take around 3 hours to complete.
bash artifact/full_exp.sh artifact/exp_one_host.csv
Print results:
python artifact/show_result_full_exp.py
G.6 Evaluation and expected result
The result should also roughly match the 2 rows in Table 5. Comparison of Methods on 1F1B in the paper:
  • 8GPU, SEQ LENGTH 2048
  • 8GPU, SEQ LENGTH 4096
The result shows that the throughput and peak memory of vocab-1 and vocab-2 are significantly better than baseline and redistribution. The throughput of interlaced is slightly better than vocab-1 and vocab-2. But the peak memory of interlaced is worse than our approach.
G.7 Experiment customization
User can customize the settings by changing the scripts quick_exp.sh, full_exp.sh, show_result_full_exp.py under under artifact/.